# Distributed Coordination for Heterogeneous Non-Terrestrial Networks

### v2

@@ VI Conclusion
In this paper, we first summarize the unique characteristics of different NTN platforms and analyze their impact on the implementation of NTNs. We then focus on investigating the communication challenges for heterogeneous NTNs in individual layers, along with their potential coordinated solutions. Based on the emerging distributed algorithm, we further investigate the distributed MADRL algorithms tailored for each potential coordinated solution in heterogeneous NTNs, and analyze the CTDE deployment options according to their latency requirement. Importantly, we utilize the TTS-MADDPG algorithm to solve the joint trajectory design and user scheduling optimization problem in a heterogeneous NTN composed of both T-UAV and U-UAV, which validates the effectiveness of the distributed coordination algorithm among heterogeneous NTN platforms.
Jikang Deng is currently a Ph.D. student at King Abdullah University of Science and Technology (KAUST), Jeddah, Saudi Arabia. His research interests include wireless communication, non-terrestrial networks, and AI for communication.
Hui Zhou is currently an Assistant Professor at Coventry University. His research interests include non-terrestrial networks, integrated sensing and communication, unlicensed spectrum, machine learning, and semantic communication.
Mohamed-Slim Alouini is currently the Al-Khawarizmi Distinguished Professor of Electrical and Computer Engineering at King Abdullah University of Science and Technology (KAUST) and the holder of the UNESCO Chair on Education to Connect the Unconnected. He is a Fellow of the IEEE and OPTICA, and his research interests include the modeling, design, and performance analysis of wireless communication systems.

### v3

@@ VI Conclusions and Future Work
In this paper, we summarize the unique characteristics of different NTN platforms and analyze their impact on the implementations. We then focus on investigating the communication challenges for heterogeneous NTNs in individual layers, along with their potential coordinated solutions. We further investigate the distributed MADRL algorithms tailored for each potential coordinated solution in heterogeneous NTNs, and analyze the CTDE deployment options according to their latency requirement. Importantly, we utilize the TTS-MADDPG algorithm to solve the joint trajectory design and user scheduling optimization problem in a heterogeneous NTN composed of both T-UAV and U-UAV, which validates the effectiveness of the distributed coordination solutions.
Moreover, in the future extended NTN scenario with a vast number of agents and multiple timescales, it is challenging to achieve effective distributed coordination, leading to slow and unstable convergence in the MADRL training. To address these issues, the MADRL can incorporate goal-oriented semantic communication and the graph encoders to reduce transmission overhead, computational load, and enhance the convergence. Moreover, it can also be extended to asynchronous MADRL based on the DTE framework, or further integrated with the grouped training and decentralized execution (GTDE) framework to improve scalability.